Suppose the charge on the electron and proton did not
have the same magnitude but instead differed by one part in 10^10. What would be the net charge on a 1.0 kg of the helium sample
1
Expert's answer
2019-09-24T09:08:58-0400

The total number of electrons per He molecule is equal to 2.

The number of moles in 1.0 kg of the helium sample is equal to

"\frac{1000}{4}=250"

The number of molecules of He in sample is equal to

250×6.02×1023=1.5×1025

The number of electrons in 1 kg of the helium sample is equal to

2×1.5×1025=3×1025

The total charge of those electrons is

3×1025×(-1.6×10-19)=-4.8×10-6 C

Answer

-4.8×10-6 C